# Monocular pose estimation of articulated surgical instruments in open surgery

### v1

@@ 3.1.3 Results on Real-World Surgical Data
GT RoI Synth 0.654 0.680 0.667
GT RoI Real 0.784 0.805 0.795
Pred RoI Real 0.754 0.752 0.753
@@
trained on synthetic data 0.564 0.455 0.510
tuned on real data 0.731 0.807 0.769
pose refinement 0.841 0.877 0.859
@@ 3.2 Ablation Studies
3.2.1 Impact of Synthetic Hand-Object Interactions
To investigate the impact of synthetic hand-object interactions on the performance of our pose estimation and object detection models, we conduct a series of ablation studies. We train three variants of our models on different synthetic datasets. These datasets are a sub-sample of the whole synthetic data that we generated:
@@ 3.2.1 Impact of Synthetic Hand-Object Interactions
    Synthetic data with masked tools: In this dataset, regions of the surgical tools that would have been occluded by hands are masked out to simulate occlusions without explicitly modeling the hands.
@@
Fig. 8: Example from hand ablation data: image with synthetic hands (left), image with masked tools (middle), image without hands (right).
@@
Without Hands 0.483 0.621 0.552
Masked Tools 0.574 0.727 0.651
With Hands 0.577 0.758 0.668
@@
For PnP-RANSAC, we utilize the Progressive-x algorithm [6] in conjunction with Graph-cut RANSAC [5]. Hereafter, we refer to this combination as PnP-RANSAC. Progressive-x is a multi-model fitting algorithm that progressively fits model instances to data points, taking into account both the fit quality and the spatial coherence of the points. Graph-cut RANSAC is a variant of RANSAC that enhances the efficiency and robustness of model fitting through graph-based segmentation. It employs a graph-based representation of the data points and uses graph-cut optimization to identify the optimal set of inliers for each model hypothesis.
@@
Table 4: Comparison of Patch-PnP and Prog-x with Graph-cut RANSAC for pose estimation inference.
Strategy Needle-holder Tweezers Mean
Patch-PnP 0.567 0.803 0.685
Prog-x 0.671 0.779 0.725
@@
Table 5: Comparison of Patch-PnP and Prog-x with Graph-cut RANSAC for pose estimation inference.
Strategy Needle-holder Tweezers Mean
Patch-PnP 0.622 0.839 0.731
Prog-x 0.784 0.805 0.795
The results show that using PnP-RANSAC overall achieves better results than Patch-PnP, though on the Tweezers, Patch-PnP outperformed PnP-RANSAC.
4 Disscusion and Conclusions
@@
While our approach shows promise, we acknowledge certain limitations that present opportunities for future research. Our current evaluation method focuses on the 2D reprojection of estimated poses, which does not fully capture the depth information of the instruments. Future work could incorporate depth information in the evaluation process to enhance 3D position accuracy. Additionally, our study was limited to two surgical tools and one site, and expanding the synthetic dataset to include a wider variety of instruments would further demonstrate the generalizability of our approach.
Looking ahead, this work opens up several directions for future research in surgical instrument pose estimation. Future studies could explore a wider range of surgical instruments and environments to further test and refine these techniques. The potential for improving both pose estimation and object detection accuracy through our iterative approach also merits further investigation.
@@
In conclusion, our study demonstrates a novel approach to pose estimation of articulated surgical instruments in open surgery using synthetic data and domain adaptation techniques. This work contributes to the ongoing efforts in computer-assisted surgery, offering a potential pathway to overcome the challenge of limited annotated real-world data. As research in this field progresses, we hope that this and similar approaches will continue to be refined, ultimately supporting the development of more effective tools for surgical assistance and training.

### v2

@@ 4.1.4 Results on Real-World Surgical Data
GT RoI Synth 0.65 0.68 0.67
GT RoI Real 0.77 0.82 0.79
Pred RoI Real 0.74 0.67 0.70
@@
trained on synthetic data 0.56 0.45 0.51
tuned on real data 0.69 0.75 0.72
pose refinement 0.71 0.76 0.74
@@ 4.2 Ablation Studies
4.2.1 Impact of Synthetic Hand-Object Occlusions
To investigate the impact of synthetic hand-object occlusions on the performance of our pose estimation and object detection models, we conduct a series of ablation studies. We train three variants of our models on different synthetic datasets. These datasets are a sub-sample of the whole synthetic data that we generated:
@@ 4.2.1 Impact of Synthetic Hand-Object Occlusions
    Synthetic data with masked tools: In this dataset, regions of the surgical tools that would have been occluded by hands were masked out using a rotating 2D bounding box that encapsulates the hand. This straightforward approach removes occluded tool regions without requiring complex hand modeling, avoiding potential biases while maintaining realistic tool occlusion patterns.
@@
Fig. 9: Examples from the synthetic hand ablation dataset illustrating occlusion variations used in Section 4.2.1. From left to right: image with synthetic hands, image without hands, and image with hands removed but tools masked at occluded regions. Rotated bounding boxes are overlaid to highlight the tool regions that are removed in the masking scenario. Additional examples are provided in the appendix.
@@
Without Hands 0.48 0.62 0.55
Masked Tools 0.56 0.73 0.64
With Hands 0.58 0.76 0.67
@@
For PnP-RANSAC, we utilize the Progressive-x algorithm [barath2019progressive] in conjunction with Graph-cut RANSAC [barath2018graph]. Hereafter, we refer to this combination as PnP-RANSAC. Progressive-x is a multi-model fitting algorithm that progressively fits model instances to data points, taking into account both the fit quality and the spatial coherence of the points. Graph-cut RANSAC is a variant of RANSAC that enhances the efficiency and robustness of model fitting through graph-based segmentation. It employs a graph-based representation of the data points and uses graph-cut optimization to identify the optimal set of inliers for each model hypothesis.
@@
Table 4: Training strategy comparison of Patch-PnP
Training Strategy Needle-holder Tweezers Mean
Without Patch-PnP 0.57 0.80 0.68
With Patch-PnP 0.67 0.78 0.72
@@
Table 5: Comparison of Patch-PnP and Prog-x for pose estimation inference.
Inference Strategy Needle-holder Tweezers Mean
Patch-PnP 0.62 0.84 0.73
Prog-x 0.78 0.80 0.79
5 Discussion
@@
5.1 Results Overview
Table 1 demonstrates the effectiveness of our approach in achieving robust pose estimation for real-world surgical tools. Our iterative domain adaptation pipeline improved the pose estimation Average Precision (AP) from 0.67 (synthetic-only training) to 0.79 when refined on real-world data with ground truth bounding boxes. When evaluated using predicted bounding boxes, the AP remains competitive at 0.70.
Similarly, object detection performance improved significantly through domain adaptation. As shown in Table 2, tuning on real data raised the detection AP from 0.51 to 0.72. Pose refinement further improved AP to 0.74.
5.2 Occlusion Ablation Results
Our occlusion ablation study (Section 4.2.1) highlighted the importance of incorporating hand-object interactions in the synthetic dataset. Training with realistic synthetic hand occlusions resulted in the highest pose estimation performance (mean AP: 0.67). These findings suggest that modeling hand occlusions is beneficial. However, we also found that simpler occlusion strategies, such as masking occluded regions without modeling hands, still yielded reasonable results, offering a practical alternative when detailed hand modeling is infeasible.
5.3 Patch-PnP vs. PnP-RANSAC Ablation Results
The ablation study comparing Patch-PnP and PnP-RANSAC (Section 4.2.2) revealed insights into their respective strengths. When evaluating inference strategies, PnP-RANSAC achieved a higher mean pose estimation AP (0.79) compared to Patch-PnP (0.73), highlighting its superior performance in leveraging robust model fitting during testing. When used as a training module, the inclusion of Patch-PnP in an end-to-end pipeline resulted in improved model optimization, yielding a mean AP of 0.72 compared to 0.68 when training without it. These results suggest that while PnP-RANSAC excels in final inference robustness, Patch-PnP contributes to improving feature learning during training, making it a valuable component for network optimization.
5.4 Benchmarking Against Prior Work
Due to the unique nature of our task, directly benchmarking against prior work proves challenging. Although laparoscopic approaches primarily address cylindrical, minimally invasive tools, existing open surgery methods (e.g., [hein2021towards, hein2023next]) focus on either joint hand and tool tracking or multi-view setups. These differences highlight the need for dedicated benchmarks and tailored methodologies in open surgery contexts. Although direct comparisons with these methods were not included in this study, future work that incorporates such comparative analyses would provide valuable insights into the relative strengths and limitations of our framework.
6 Future Work
@@
While our approach shows promise, we acknowledge certain limitations that present opportunities for future research. Our current evaluation method focuses on the 2D reprojection of estimated poses, which does not fully capture the depth information of the instruments. Future work could incorporate depth information in the evaluation process to enhance 3D position accuracy.
Additionally, our study was limited to two surgical tools and one site, and expanding the synthetic dataset to include a wider variety of instruments would further demonstrate the generalizability of our approach.
Furthermore, our approach utilized several criteria to assess the quality of the pseudo-labels. Although these criteria effectively facilitated synthetic-to-real domain adaptation, a comprehensive hyperparameter search to evaluate their sensitivity and impact on downstream performance remains unexplored. Future work should investigate how these parameters influence the framework’s robustness across diverse domains.
7 Conclusions
In conclusion, this study demonstrates a novel approach to pose estimation of articulated surgical instruments in open surgery. Our approach involved creating a diverse synthetic dataset of surgical instruments with various articulation angles, enabling initial model training. These models were then refined through synthetic to real domain adaptation techniques, resulting in robust predictions of 6D pose, instrument class, and articulation angle in real surgical scenarios. The results demonstrate our method’s effectiveness in addressing unique surgical environment challenges, such as limited real-world data, complex instrument geometry, and occlusions. Notably, our approach shows promise not only in pose estimation but also in improving object detection accuracy through an iterative refinement process.
@@ 7 Conclusions
This work contributes to ongoing efforts in computer-assisted surgery, offering a potential pathway to overcome the challenge of limited annotated real-world data. As research in this field progresses, we hope that this and similar approaches will continue to be refined, ultimately supporting the development of more effective tools for surgical assistance and training.